# Existence of Friedrich-Wintgen Bound States in the Continuum: Cavity with a Thin Waveguide Opening

### v1

@@ 5.2 Continuity in $\delta$ and Lipschitz continuity in $\mu$ for $f_{0}$ and $f_{1}$
In the following analysis, we establish the continuity of the functions $f_{0}(\delta,\mu,h)$ and $f_{1}(\delta,\mu,h)$ in the variable $\delta$, as well as their Lipschitz continuity in $\mu$ with a Lipschitz constant less than $1$.
@@
Under Condition A.1, $\mathbf{T}(\delta,\mu)$ is continues in $\delta$ and $\mu$ for $\delta$ in a sufficiently small interval and $\mu$ satisfying (38). Moreover, there is a constant $C_{2}$ irrelevant to $\delta$ that
(92) $\|\mathbf{T}(\delta,\mu_{1})-\mathbf{T}(\delta,\mu_{2})\|_{\ell^{2}\times\ell^{2}}\leq{C_{2}}|\mu_{1}-\mu_{2}|.$
@@
Given $m\geq{M}$, each element in $\mathbf{T}(\delta_{1},\mu)-\mathbf{T}(\delta_{2},\mu)$ satisfies
$\left|\frac{\sqrt{1+\lambda_{m}(\delta_{1})}}{\sqrt{\lambda_{m}(\delta_{1})-\mu}}-\frac{\sqrt{1+\lambda_{m}(\delta_{2})}}{\sqrt{\lambda_{m}(\delta_{2})-\mu}}\right|$
$=$ $\left|\left(\sqrt{1+\frac{\mu+1}{\lambda_{m}(\delta_{1})-\mu}}+\sqrt{1+\frac{\mu+1}{\lambda_{m}(\delta_{2})-\mu}}\right)^{-1}\left(\frac{\mu+1}{\lambda_{m}(\delta_{1})-\mu}-\frac{\mu+1}{\lambda_{m}(\delta_{2})-\mu}\right)\right|$
$\leq$ $\frac{1}{2}(\lambda_{M}(0)+1)\left(1+\frac{\lambda_{M}(0)+1}{\epsilon}\right)^{2}\left(e^{C_{0}|\delta_{1}-\delta_{2}|}-1\right),$
which tends to 00 uniformly. Then the continuity of $\mathbf{T}(\delta,\mu)$ with respect to $\delta$ follows. Similarly, each element in $\mathbf{T}(\delta,\mu_{1})-\mathbf{T}(\delta,\mu_{2})$ satisfies
$\left|\frac{\sqrt{1+\lambda_{m}(\delta)}}{\sqrt{\lambda_{m}(\delta)-\mu_{1}}}-\frac{\sqrt{1+\lambda_{m}(\delta)}}{\sqrt{\lambda_{m}(\delta)-\mu_{2}}}\right|$
$=$ $\sqrt{1+\frac{\mu_{1}+1}{\lambda_{m}(\delta)-\mu_{1}}}\left|\frac{\mu_{1}-\mu_{2}}{\sqrt{\lambda_{m}(\delta)-\mu_{2}}(\sqrt{\lambda_{m}(\delta)-\mu_{1}}+\sqrt{\lambda_{m}(\delta)-\mu_{2}})}\right|$
$\leq$ $\frac{1}{2\epsilon}\sqrt{1+\frac{\lambda_{M}(0)+1}{\epsilon}}|\mu_{1}-\mu_{2}|,$
@@
To study the continuity of the operator $(\mathbf{I}-\mathbf{B})^{-1}$ in $\ell^{2}$ with respect to $\delta$, we say an operator $\mathbf{G}(\delta)$ in $\ell^{2}$ is strongly continuous on $[\delta_{a},\delta_{b}]\subset\mathbb{R}$ if for every $\mathbf{d}\in\ell^{2}$
(93) $\left\|\mathbf{G}(\delta_{1})\mathbf{d}-\mathbf{G}(\delta_{2})\mathbf{d}\right\|_{\ell^{2}}\to 0\ \text{as}\ \delta_{1}\to\delta_{2},\ \text{for}\ \delta_{1},\delta_{2}\in[\delta_{a},\delta_{b}],$
as defined in [10, p. 152].
@@
Let $\mathbf{G}(\delta)$ denote an invertible, self-adjoint operator in $\ell^{2}$ that is strongly continuous with respect to $\delta\in[\delta_{a},\delta_{b}]$, where $\delta_{a},\delta_{b}\in\mathbb{R}$. If its inverse $\mathbf{G}^{-1}(\delta)$ is uniformly bounded on $[\delta_{a},\delta_{b}]$, then $\mathbf{G}^{-1}(\delta)$ is also strongly continuous.
@@
(94) $(\mathbf{v}(\delta_{1}),\mathbf{G}(\delta_{1})\mathbf{q}-\mathbf{G}(\delta_{2})\mathbf{q})_{\ell^{2}}\to 0\ \text{as}\ \delta_{1}\to\delta_{2}.$
Define $\widetilde{\mathbf{q}}:=\mathbf{G}(\delta_{2})\mathbf{q}$ and $\mathbf{v}(\delta_{1}):=\mathbf{G}^{-1}(\delta_{1})(\mathbf{G}^{-1}(\delta_{2})\widetilde{\mathbf{q}}-\mathbf{G}^{-1}(\delta_{1})\widetilde{\mathbf{q}})$. Substituting into (94) and using the self-adjointness of $\mathbf{G}(\delta_{1})$ and uniform boundedness of $\mathbf{G}^{-1}(\delta_{1})$, we obtain
(95) $(\mathbf{G}^{-1}(\delta_{2})\widetilde{\mathbf{q}}-\mathbf{G}^{-1}(\delta_{1})\widetilde{\mathbf{q}},\mathbf{G}^{-1}(\delta_{2})\widetilde{\mathbf{q}}-\mathbf{G}^{-1}(\delta_{1})\widetilde{\mathbf{q}})_{\ell^{2}}\to 0\ \text{as}\ \delta_{1}\to\delta_{2},$
@@
Let $\mathbf{H}(\delta,\mu)$ denote an operator in $\ell^{2}$ defined by
@@
Under Condition A.1 and the assumption $h\ll 1$, the operator $\mathbf{H}(\delta,\mu)$ is strongly continuous in $\delta$ for $\delta$ in a sufficiently small interval and $\mu$ satisfying (38). Furthermore, we have
(98) $\|\mathbf{H}(\delta,\mu_{1})-\mathbf{H}(\delta,\mu_{2})\|_{\ell^{2}\times\ell^{2}}\leq{C_{3}}|\mu_{1}-\mu_{2}|,$
@@
We consider the operator $\widetilde{\mathbf{H}}(\delta,\mu)$ defined as
@@
where
@@
Accordingly, $\mathbf{H}(\delta,\mu)$ serves as a restriction of $\widetilde{\mathbf{H}}(\delta,\mu)$. Let $\ell^{2}_{c}$ denote a subspace of $\ell^{2}$ defined by
(101) $\ell^{2}_{c}:=\left\{\{d_{m}\}\in\ell^{2}:\exists{m_{0}\in\mathbb{N}}\ s.t.\ d_{m}=0\ \text{if}\ m{\geq}m_{0}\right\},$
@@
which is dense in $\ell^{2}$. Given $\{\widetilde{d}_{0,m}\}\in\ell^{2}_{c}$, we first show the continuity of $\mathbf{d}(\delta):=\widetilde{\mathbf{H}}(\delta,\mu)\{\widetilde{d}_{0,m}\}$ in $\ell^{2}$. For any $\{\widetilde{d}_{1,m}\}\in\ell^{2}_{c}$, the inner product $(\mathbf{d}(\delta_{1}),\{\widetilde{d}_{1,m}\})_{\ell^{2}}$ converges to $(\mathbf{d}(\delta_{2}),\{\widetilde{d}_{1,m}\})_{\ell^{2}}$ as $\delta_{1}\to\delta_{2}$ by Lemma 3.1. Hence, the Banach-Steinhaus Theorem (cf. [20]) implies the weak convergence of $\mathbf{d}(\delta_{1})$ to $\mathbf{d}(\delta_{2})$. Let $\{d_{0,m}\}\in\ell^{2}$ with $\|\{d_{0,m}\}\|_{\ell^{2}}=1$ such that
$\|\mathbf{d}(\delta_{2})\|_{\ell^{2}}$ $=(\mathbf{d}(\delta_{2}),\{d_{0,m}\})_{\ell^{2}}$
@@
(102) $=\sum_{j=1}^{\infty}{\bf i}\alpha_{j}(\mu)(\phi_{j},\sum_{m=0}^{\infty}\frac{\widetilde{d}_{0,m}\psi_{m}(\cdot,\delta_{2})}{\sqrt{1+\lambda_{m}(\delta_{2})}})_{\Gamma_{h}}(\sum_{m=0}^{\infty}\frac{d_{0,m}\psi_{m}(\cdot,\delta_{2})}{\sqrt{1+\lambda_{m}(\delta_{2})}},\phi_{j})_{\Gamma_{h}}.$
@@
Since $\{{\psi_{m}(\cdot,\delta)}/{\sqrt{1+\lambda_{m}(\delta)}}\}$ forms a complete orthonormal basis in $H^{1}(\Omega_{\mathrm{in}})$, there exists $\{d_{1,m}\}\in\ell^{2}$ such that $\|\{d_{1,m}\}\|_{\ell^{2}}=1$ and
@@
(103) $\sum_{m=0}^{\infty}\frac{d_{0,m}\psi_{m}(\cdot,\delta_{2})}{\sqrt{1+\lambda_{m}(\delta_{2})}}\Bigg{|}_{\Gamma_{h}}=\sum_{m=0}^{\infty}\frac{d_{1,m}\psi_{m}(\cdot,\delta_{1})}{\sqrt{1+\lambda_{m}(\delta_{1})}}\Bigg{|}_{\Gamma_{h}}.$
Hence, we have
$\|\mathbf{d}(\delta_{2})\|_{\ell^{2}}-\|\mathbf{d}(\delta_{1})\|_{\ell^{2}}$ $\leq\|\mathbf{d}(\delta_{2})\|_{\ell^{2}}-(\mathbf{d}({\delta_{1}}),\{d_{1,m}\})_{\ell^{2}}$
@@
(104) ${\leq}n_{1}^{2}$ $\left\|\frac{1}{n(\cdot,\delta_{1})}\nabla\left(\sum_{m=0}^{\infty}\frac{\widetilde{d}_{0,m}\psi_{m}(\cdot,\delta_{2})}{\sqrt{1+\lambda_{m}(\delta_{2})}}-\sum_{m=0}^{\infty}\frac{\widetilde{d}_{0,m}\psi_{m}(\cdot,\delta_{1})}{\sqrt{1+\lambda_{m}(\delta_{1})}}\right)\right\|_{[L^{2}(R_{h})]^{2}}$
by Lemma 4.3. Furthermore, (104) becomes infinitesimally small as $\delta_{1}\to\delta_{2}$ by Lemma 3.1, since $\{\widetilde{d}_{0,m}\}\in\ell^{2}_{c}$. Exchanging $\delta_{1}$ and $\delta_{2}$ in (104) yields
(105) $\|\mathbf{d}(\delta_{1})\|_{\ell^{2}}\to\|\mathbf{d}(\delta_{2})\|_{\ell^{2}},\ \text{as}\ \delta_{1}\to\delta_{2},$
@@
which implies the strong convergence of $\mathbf{d}(\delta_{1})$ to $\mathbf{d}(\delta_{2})$ (cf. Exercise 2.11 in [18]). Therefore $\widetilde{\mathbf{H}}(\delta,\mu)$ is strongly continuous by the Banach-Steinhaus Theorem.
To prove (98), consider the inner product for given $\{d_{0,m}\},\{d_{1,m}\}\in\ell^{2}$:
$\left|((\widetilde{\mathbf{H}}(\delta,\mu_{1})-\widetilde{\mathbf{H}}(\delta,\mu_{2}))\{d_{0,m}\},\{d_{1,m}\})_{\ell^{2}}\right|$
@@
$=$ $\sum_{j=1}^{\infty}{\bf i}\left|(\alpha_{j}(\mu_{1})-\alpha_{j}(\mu_{2}))\right|(\phi_{j},\sum_{m=0}^{\infty}\frac{d_{0,m}\psi_{m}(\cdot,\delta)}{\sqrt{1+\lambda_{m}(\delta)}})_{\Gamma_{h}}(\phi_{j},\sum_{m=0}^{\infty}\frac{d_{1,m}\psi_{m}(\cdot,\delta)}{\sqrt{1+\lambda_{m}(\delta)}})_{\Gamma_{h}}$
(106) $\leq$ $C_{3}|\mu_{1}-\mu_{2}|\|\{d_{0,m}\}\|_{\ell^{2}}\|\{d_{1,m}\}\|_{\ell^{2}},$
where we have used the same techniques as in Lemma 4.3 and the assumption $h\ll 1$.
@@
Under Conditions A.1, A.2, and the assumption $h\ll 1$, the functions $f_{0}(\delta,\mu,h)$ and $f_{1}(\delta,\mu,h)$ are continuous in $\delta$ and Lipschitz continuous in $\mu$ for $\delta$ in a sufficiently small interval and $\mu$ satisfying (38). Furthermore, the following Lipschitz bounds hold:
(107) $|f_{0}(\delta,\mu_{1},h)-f_{0}(\delta,\mu_{2},h)|\leq$ ${C_{f_{0}}}|\mu_{1}-\mu_{2}|,$
(108) $|f_{1}(\delta,\mu_{1},h)-f_{1}(\delta,\mu_{2},h)|\leq$ ${C_{f_{1}}}|\mu_{1}-\mu_{2}|,$
where $C_{f_{0}}<1$ and $C_{f_{1}}<1$ are two constants independent of $\delta$.

### v2

@@ 5.2 Continuity in $\delta$ and Lipschitz continuity in $\mu$ for $f_{0}$ and $f_{1}$
In the following analysis, we establish the continuity of the functions $f_{0}(\delta,\mu,h)$ and $f_{1}(\delta,\mu,h)$ in $\delta$, as well as their Lipschitz continuity in $\mu$ with a Lipschitz constant less than $1$.
@@
Under Condition A.1, $\mathbf{T}(\delta,\mu)$ is continues in $\delta$ and $\mu$ for $\delta\in[-\delta_{1},\delta_{1}]$ and $\mu\in[\mu_{l},\mu_{r}]$. Moreover, there is a constant $C_{2}$ irrelevant to $\delta$ such that
(95) $\|\mathbf{T}(\delta,\mu_{a})-\mathbf{T}(\delta,\mu_{b})\|_{\ell^{2}\to\ell^{2}}\leq{C_{2}}|\mu_{a}-\mu_{b}|,\ \text{for}\ \mu_{a},\mu_{b}\in[\mu_{l},\mu_{r}].$
@@
Given $m\geq{M}$, each element in $\mathbf{T}(\delta_{a},\mu)-\mathbf{T}(\delta_{b},\mu)$ satisfies
$\left|\frac{\sqrt{1+\lambda_{m}(\delta_{a})}}{\sqrt{\lambda_{m}(\delta_{a})-\mu}}-\frac{\sqrt{1+\lambda_{m}(\delta_{b})}}{\sqrt{\lambda_{m}(\delta_{b})-\mu}}\right|$
$=$ $\left|\left(\sqrt{1+\frac{\mu+1}{\lambda_{m}(\delta_{a})-\mu}}+\sqrt{1+\frac{\mu+1}{\lambda_{m}(\delta_{b})-\mu}}\right)^{-1}\left(\frac{\mu+1}{\lambda_{m}(\delta_{a})-\mu}-\frac{\mu+1}{\lambda_{m}(\delta_{b})-\mu}\right)\right|$
$\leq$ $\frac{1}{2}(\lambda_{M}(0)+1)\left(1+\frac{\lambda_{M}(0)+1}{\epsilon}\right)^{2}\left(e^{C_{0}|\delta_{a}-\delta_{b}|}-1\right),$
which tends to 0 uniformly. Then the continuity of $\mathbf{T}(\delta,\mu)$ with respect to $\delta$ follows. Similarly, each element in $\mathbf{T}(\delta,\mu_{a})-\mathbf{T}(\delta,\mu_{b})$ satisfies
$\left|\frac{\sqrt{1+\lambda_{m}(\delta)}}{\sqrt{\lambda_{m}(\delta)-\mu_{a}}}-\frac{\sqrt{1+\lambda_{m}(\delta)}}{\sqrt{\lambda_{m}(\delta)-\mu_{b}}}\right|$
$=$ $\sqrt{1+\frac{\mu_{a}+1}{\lambda_{m}(\delta)-\mu_{a}}}\left|\frac{\mu_{a}-\mu_{b}}{\sqrt{\lambda_{m}(\delta)-\mu_{b}}(\sqrt{\lambda_{m}(\delta)-\mu_{a}}+\sqrt{\lambda_{m}(\delta)-\mu_{b}})}\right|$
$\leq$ $\frac{1}{2\epsilon}\sqrt{1+\frac{\lambda_{M}(0)+1}{\epsilon}}|\mu_{a}-\mu_{b}|,$
@@
To study the continuity of the operator $(\mathbf{I}-\mathbf{B})^{-1}$ in $\ell^{2}$ with respect to $\delta$, we say an operator $\mathbf{G}(\delta)$ in $\ell^{2}$ is strongly continuous on $\delta_{0},\delta_{0}]$ if for every $\mathbf{d}\in\ell^{2}$
(96) $\left\|\mathbf{G}(\delta_{a})\mathbf{d}-\mathbf{G}(\delta_{b})\mathbf{d}\right\|_{\ell^{2}}\to 0\ \text{as}\ \delta_{a}\to\delta_{b},\ \text{for}\ \delta_{a},\delta_{b}\in[-\delta_{0},\delta_{0}],$
as defined in [11, p. 152].
@@
Let $\mathbf{G}(\delta)$ denote an invertible, self-adjoint operator in $\ell^{2}$ that is strongly continuous with respect to $\delta\in[-\delta_{0},\delta_{0}]$. If its inverse $\mathbf{G}^{-1}(\delta)$ is uniformly bounded on $\delta_{0},\delta_{0}]$, then $\mathbf{G}^{-1}(\delta)$ is also strongly continuous.
@@
(97) $(\mathbf{v}(\delta_{a}),\mathbf{G}(\delta_{a})\mathbf{q}-\mathbf{G}(\delta_{b})\mathbf{q})_{\ell^{2}}\to 0\ \text{as}\ \delta_{a}\to\delta_{b}.$
[…]
(98) $(\mathbf{G}^{-1}(\delta_{b})\widetilde{\mathbf{q}}-\mathbf{G}^{-1}(\delta_{a})\widetilde{\mathbf{q}},\mathbf{G}^{-1}(\delta_{b})\widetilde{\mathbf{q}}-\mathbf{G}^{-1}(\delta_{a})\widetilde{\mathbf{q}})_{\ell^{2}}\to 0\ \text{as}\ \delta_{a}\to\delta_{b},$
@@
Let $\mathbf{H}(\delta,\mu)$ be an operator in $\ell^{2}$ defined by
@@
Under Condition A.1 and for $h\in(0,h_{0})$, the operator $\mathbf{H}(\delta,\mu)$ is strongly continuous in $\delta$ for $\delta\in[-\delta_{2},\delta_{2}]$ and $\mu\in[\mu_{l},\mu_{r}]$. Moreover, it satisfies the Lipschitz estimate
(101) $\|\mathbf{H}(\delta,\mu_{a})-\mathbf{H}(\delta,\mu_{b})\|_{\ell^{2}\to\ell^{2}}\leq{C_{3}}|\mu_{a}-\mu_{b}|\ \text{for}\ \mu_{a},\mu_{b}\in[\mu_{l},\mu_{r}],$
@@
Consider the operator
@@
where the vectors $\widetilde{\mathbf{w}}_{j}(\delta)$ are defined by
@@
The operator $\mathbf{H}(\delta,\mu)$ is then a restriction of $\widetilde{\mathbf{H}}(\delta,\mu)$. Let $\ell^{2}_{c}$ be a dense subspace of $\ell^{2}$ consisting of sequences with finite support:
(104) $\ell^{2}_{c}:=\left\{\{\widetilde{d}_{m}\}\in\ell^{2}:\exists{m_{0}\in\mathbb{N}}\ s.t.\ d_{m}=0\ \text{if}\ m{\geq}m_{0}\right\}.$
@@
Letting $\{\widetilde{d}_{0,m}\}\in\ell^{2}_{c}$, we prove the continuity of $\mathbf{d}(\delta):=\widetilde{\mathbf{H}}(\delta,\mu)\{\widetilde{d}_{0,m}\}$ in $\ell^{2}$. For any $\{\widetilde{d}_{1,m}\}\in\ell^{2}_{c}$, Lemma 3.1 implies that the inner product $(\mathbf{d}(\delta_{a}),\{\widetilde{d}_{1,m}\})_{\ell^{2}}$ converges to $(\mathbf{d}(\delta_{b}),\{\widetilde{d}_{1,m}\})_{\ell^{2}}$ as $\delta_{a}\to\delta_{b}$ for $\delta_{a},\delta_{b}\in[-\delta_{2},\delta_{2}]$. By the Banach-Steinhaus Theorem (cf. Theorem 2.7 in [23]), it follows that $\mathbf{d}(\delta_{a})$ converges weakly to $\mathbf{d}(\delta_{b})$. Now, let $\{d_{0,m}\}\in\ell^{2}$ with $\|\{d_{0,m}\}\|_{\ell^{2}}=1$ such that
$\|\mathbf{d}(\delta_{b})\|_{\ell^{2}}$ $=(\mathbf{d}(\delta_{b}),\{d_{0,m}\})_{\ell^{2}}$
@@
(105) $=\sum_{j=1}^{\infty}{\bf i}\alpha_{j}(\mu)(\phi_{j},\sum_{m=0}^{\infty}\frac{\widetilde{d}_{0,m}\psi_{m}(\cdot,\delta_{b})}{\sqrt{1+\lambda_{m}(\delta_{b})}})_{\Gamma_{h}}(\sum_{m=0}^{\infty}\frac{d_{0,m}\psi_{m}(\cdot,\delta_{b})}{\sqrt{1+\lambda_{m}(\delta_{b})}},\phi_{j})_{\Gamma_{h}}.$
@@
Since $\{{\psi_{m}(\cdot,\delta)}/{\sqrt{1+\lambda_{m}(\delta)}}\}$ forms a complete orthonormal basis in $H^{1}(\Omega_{\mathrm{in}})$, there exists a sequence $\{d_{1,m}\}\in\ell^{2}$ with $\|\{d_{1,m}\}\|_{\ell^{2}}=1$ such that
@@
(106) $\sum_{m=0}^{\infty}\frac{d_{0,m}\psi_{m}(\cdot,\delta_{b})}{\sqrt{1+\lambda_{m}(\delta_{b})}}\Bigg|_{\Gamma_{h}}=\sum_{m=0}^{\infty}\frac{d_{1,m}\psi_{m}(\cdot,\delta_{a})}{\sqrt{1+\lambda_{m}(\delta_{a})}}\Bigg|_{\Gamma_{h}}.$
Consequently, we have the estimate
$\|\mathbf{d}(\delta_{b})\|_{\ell^{2}}-\|\mathbf{d}(\delta_{a})\|_{\ell^{2}}$ $\leq\|\mathbf{d}(\delta_{b})\|_{\ell^{2}}-(\mathbf{d}({\delta_{a}}),\{d_{1,m}\})_{\ell^{2}}$
@@
(107) ${\leq}n_{1}^{2}$ $\left\|\frac{1}{n(\cdot,\delta_{a})}\nabla\left(\sum_{m=0}^{\infty}\frac{\widetilde{d}_{0,m}\psi_{m}(\cdot,\delta_{b})}{\sqrt{1+\lambda_{m}(\delta_{b})}}-\sum_{m=0}^{\infty}\frac{\widetilde{d}_{0,m}\psi_{m}(\cdot,\delta_{a})}{\sqrt{1+\lambda_{m}(\delta_{a})}}\right)\right\|_{[L^{2}(R_{h})]^{2}}$
where the second inequality follows from Lemma 4.3. Since $\{\widetilde{d}_{0,m}\}\in\ell^{2}_{c}$, Lemma 3.1 implies that the right-hand side of (107) vanishes as $\delta_{a}\to\delta_{b}$. Exchanging $\delta_{a}$ and $\delta_{b}$ shows that
(108) $\|\mathbf{d}(\delta_{a})\|_{\ell^{2}}\to\|\mathbf{d}(\delta_{b})\|_{\ell^{2}},\ \text{as}\ \delta_{a}\to\delta_{b}.$
@@
This convergence of norms, together with the established weak convergence, implies the $\ell^{2}$ convergence of $\mathbf{d}(\delta_{a})$ to $\mathbf{d}(\delta_{b})$ (cf. Exercise 2.11 in [20]). Therefore $\widetilde{\mathbf{H}}(\delta,\mu)$ is strongly continuous by the Banach-Steinhaus Theorem.
We now prove the Lipschitz continuity in $\mu$, i.e., (101). For any $\{d_{0,m}\},\{d_{1,m}\}\in\ell^{2}$, we have
$\left|((\widetilde{\mathbf{H}}(\delta,\mu_{a})-\widetilde{\mathbf{H}}(\delta,\mu_{b}))\{d_{0,m}\},\{d_{1,m}\})_{\ell^{2}}\right|$
@@
$=$ $\sum_{j=1}^{\infty}\left|{\bf i}(\alpha_{j}(\mu_{a})-\alpha_{j}(\mu_{b}))\right|(\phi_{j},\sum_{m=0}^{\infty}\frac{d_{0,m}\psi_{m}(\cdot,\delta)}{\sqrt{1+\lambda_{m}(\delta)}})_{\Gamma_{h}}(\sum_{m=0}^{\infty}\frac{d_{1,m}\psi_{m}(\cdot,\delta)}{\sqrt{1+\lambda_{m}(\delta)}},\phi_{j})_{\Gamma_{h}}$
(109) $\leq$ $\frac{1}{2\epsilon^{2}}|\mu_{a}-\mu_{b}|\|\{d_{0,m}\}\|_{\ell^{2}}\|\{d_{1,m}\}\|_{\ell^{2}},\ \text{for}\ \mu_{a},\mu_{b}\in[\mu_{l},\mu_{r}].$
The inequality relies on Lemma 4.3 and the bound
(110) $|{\bf i}(\alpha_{j}(\mu_{a})-\alpha_{j}(\mu_{b}))|\leq\frac{{\bf i}\alpha_{j}(\mu_{b})|\mu_{b}-\mu_{a}|}{({\bf i}\alpha_{j}(\mu_{a})+{\bf i}\alpha_{j}(\mu_{b})){\bf i}\alpha_{j}(\mu_{b})}\leq\frac{{\bf i}\alpha_{j}(\mu_{b})|\mu_{b}-\mu_{a}|}{2\epsilon^{2}}.$
@@
Under Conditions A.1, A.2, there exists $h_{3}>0$ such that for all $h\in(0,h_{3})$, $\delta\in[-\delta_{3},\delta_{3}]$ and $\mu\in[\mu_{l},\mu_{r}]$, the functions $f_{0}(\delta,\mu,h)$ and $f_{1}(\delta,\mu,h)$ are continuous in $\delta$ and Lipschitz continuous in $\mu$. Moreover, the following Lipschitz estimates hold for all $\mu_{a},\mu_{b}\in[\mu_{l},\mu_{r}]$:
(111) $|f_{0}(\delta,\mu_{a},h)-f_{0}(\delta,\mu_{b},h)|\leq$ ${C_{f_{0}}}|\mu_{a}-\mu_{b}|,$
(112) $|f_{1}(\delta,\mu_{a},h)-f_{1}(\delta,\mu_{b},h)|\leq$ ${C_{f_{1}}}|\mu_{a}-\mu_{b}|,$
where the constants $C_{f_{0}}<1$ and $C_{f_{1}}<1$ are independent of $\delta$ and $h$.